8) Is the point (-5, 3) on the circle defined by (x - 2)2 + (y + 1)2 = 64? Justify your answer! A) No B) Yes
Answer #1

equation of ce px - 2)² + (y -k} = 2 (ls, k) center & r radius (x - 2)2 +ly to)2 = 64 (x-2) + (y=(-1)}² = 82 center (+29-1) r